Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.20 pounds $ 16.70 per pound $ 36.74
Direct labor 1.00 hours $ 15.60 per hour $ 15.60
Variable manufacturing overhead 1.00 hours $ 9.30 per hour $ 9.30
Total standard cost per unit $ 61.64
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 587,840 $ 10,788 F $ 33,400 U
Direct labor $ 249,600 $ 3,400 U $ 15,600 U
Variable manufacturing overhead $ 148,800 $ 4,000 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many units were produced last period?

2. How many pounds of direct material were purchased and used in production?

3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

4. How many actual direct labor-hours were worked during the period?

5. What was the actual rate paid per direct labor-hour? (Round your answer to 2 decimal places.)

6. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

1. No of units produced = Total standard cost/standard cost per unit =587840/36.74 = 16000


2. Pounds of direct material purchased and used in production :

Material quantity variance = (standard quantity-actual quantity)standard price

                           -33400 = (16000*2-X)16.70

                           -33400 = (32000-x)17.60

                           -33400 = 563200-16.70X

                           16.70X = 596600

X(actual quantity purchased and used) = 35724 pounds

3. Actual cost per pound :

material price variance = (standard price-actual price)actual quantity

                      10788 = (16.70-X)35724

                      10788 = 596591-35724X

                   35724X     = 585803

X(actual cost per pound) = 16.39

4.) calculate actual direct labour hours worked :

Direct labour efficiency variance = (standard hours-actual hours)standard rate

                          -15600    = (16000-X)15.60

                                  -15600     = 249600-15.60X

                                  15.60X     = 265200

     x(actual direct labour hours)   = 17000 Hours

5. Actual rate per direct labour hour :

Labour rate variance = (standard rate-actual rate)actual hours

                  -3400     = (15.60-X)17000

                  -3400      = 265200-17000X

                17000X      = 268600

X)(actual rate per direct labour hour) = 15.8 per hour

6) actual variable overhead cost :

Actual variable overhead = standard overhead+Unfavourable variance

                                    = 148800+5300

Actual variance overhead = 154100

note: Using the numbers from #4, take the standard DLH - actual DLH to get DLH (Unfavorable). 16000-17000=-1000. Now multiply 1000 x standard overhead rate $9.30. -1000 x 9.30 = -9300. Now add -9300 to the favorable rate $4000. -9300+4000=5300.